# Impact of Sunglasses on One-to-Many Facial Identification Accuracy

### v1

@@ IV-B Effect of Sunglasses + Blur / Resolution
TABLE II: Wasserstein distance for shift from baseline caused by effects in probe image. Distributions in Figure 7. Face matcher is AdaFace.
@@
TABLE III: FPIR for one-to-many identification experiments under various effects in the probe image. As an example, with sunglasses + blur in the probe images, 4.522% of Caucasian females probe images had a rank-one false positive identification with AdaFace.
Figure 6 shows one example of probe images for six conditions: 1) Original high-quality image, 2) Sunglasses added to original, 3) $\sigma=4.6$ Gaussian blur added to original, 4) Original resized to 37x37, 5) Sunglasses and blur, and 6) Sunglasses and reduced resolution. Note that adding sunglasses to an original image may subjectively appear to result in a high-quality image, even though accuracy is degraded similar to a noticeable degree of blur or low resolution. Table I shows that when combining blur and low resolution with sunglasses, the d-prime values are almost doubled between mated distributions and almost tripled between non-mated distributions, compared with the single factor, for both face matchers and demographic groups. For example, in the AdaFace results for mated females, the d-prime values are 2.6694 when only adding sunglasses to the probes and 5.783 when combining sunglasses and blur in the probes.
@@ IV-C Analysis of Effect on Model Performance
To quantitatively analyze the false positive identifications (FPI) caused by each condition, we report the Wasserstein Distance of the score difference distribution and the FPIR. For a given probe image, a negative score difference indicates an FPI, as the highest similarity value between the probe image and an image from a different person is higher than the top matched image from the same identity. Thus, the distribution of score differences provides a measure of the FPI potential. The distributions of these differences are shown in Figure 7. A general conclusion is that each of the probe quality factors – wearing sunglasses, increased blur, lower resolution – causes the (mated - non-mated) score distribution to shift to lower values, representing increased chances for FPI. The greatest shifts naturally occur when probe images have combined blur and sunglasses, or combined lower-resolution and sunglasses.
@@
We also calculated the False Positive Identification Rate (FPIR) for each probe quality condition, using both the ArcFace and AdaFace matchers. These results in Table III show that both ArcFace and AdaFace have zero FPIR for the original and 37x37 resolution probes. AdaFace experiences an FPIR of 0.522% for sunglasses-added probes and an FPIR of 0.174% for blurred probes. But the FPIR increases sharply when sunglasses are combined with either blur or reduced resolution. For example, sunglasses + blur results in FPIR of 3.652% for females for ArcFace, and 4.522% for AdaFace. The absolute level of FPIR varies with the size of a dataset, with a larger number of persons enrolled in the gallery causing an increased FPIR [24]. So the overall low FPIR in these results does not project what can be experienced in practice for scenarios with larger galleries. However, the important general point illustrated in these results is that multiple degraded quality factors in the probe image causes a dramatic increase in the FPIR. Also, both matchers consistently show higher FPIR for females than for males. This gender-based difference illustrates another general lesson for one-to-many identification accuracy.
@@ IV-D Strategy for Recovering Accuracy Loss
Previous sections show that a probe image with sunglasses matched against a gallery of images without sunglasses results in a substantial decrease in accuracy. Primarily, the occlusion of facial features by sunglasses in the probe image obscures vital identifying information. But also, comparing images with and without sunglasses presents a mismatch in facial features, which is potentially a different factor affecting accuracy. One could imagine an approach to remove sunglasses from the probe image, or an approach to add sunglasses to the gallery images. Removing sunglasses from the probe could perhaps require less computation, as it involves the one image, but it requires hallucinating face feature content that is unseen. Adding sunglasses to the gallery images, on the other hand, involves adding a plausible instance of a specific style of occlusion to each image.
@@
Table IV The table illustrates the d-prime values of rank-one similarity distributions for probe images with sunglasses under different gallery conditions. When sunglasses are randomly added to a single image per identity in the gallery, with an average of 11.15 images per identity, the results show d-prime values of 0.5344 (mated) and 0.7142 (non-mated) for females, and 0.6911 (mated) and 0.891 (non-mated) for males. The d-prime values become more pronounced when sunglasses are added to all gallery images, as evidenced by the increased values: 1.3203 (mated) and 1.6576 (non-mated) for females, and 1.619 (mated) and 1.8357 (non-mated) for males. Notably, in both scenarios, the improvement in accuracy is more substantial for males than for females. The results demonstrate that under both gallery conditions, accuracy improves; however, adding sunglasses to all gallery images leads to a higher recovery in accuracy. The value of adding sunglasses to the gallery images to ameliorate the accuracy degradation is more obvious in the (mated - non-mated) distributions shown in Figure 8.
@@ IV-E Effect of Sunglasses Training Samples in Accuracy
Accuracy lost by adding sunglasses to face can possibly be explained by the under-representation of wearing-sunglasses samples in the training set. Thus, we start the analysis on a widely recognized training set, WebFace4M [40], which encompasses 4M images from 200K identities.
Considering the absence of dark sunglasses labels of the WebFace4M training set and off-the-shelf sunglasses classifiers, we developed a specialized dark sunglasses detector to obtain the number of wearing-sunglasses samples in WebFace4M. Specifically, we first use a facial attribute classifier [36] to get the wearing-eyeglasses images in WebFace4M. Part of these images are then manually separated to three classes: no-eyeglasses, eyeglasses, and sunglasses, for the next training step. This process results in 8,500 dark sunglasses images, 8,200 eyeglasses images, and 8,200 no-eyeglasses images. This dataset is subsequently partitioned into training, validation, and test sets, with allocations of 70%, 15%, and 15%, respectively. We then fine-tuned a ResNet101 [12], which is pre-trained on ImageNet, using this dataset. This dark sunglasses detector achieves 93.3% accuracy on the test set. After running this detector on WebFace4M, there are 115,818 (3.8%) images are predicted as having dark sunglasses, showcasing the wearing-sunglasses samples are under-represented in the training set. Hence, increasing the representation of wearing-sunglasses images in the training set is a possible way to improve the accuracy.
@@
However, adding sunglasses images to the dataset is non-trivial, so we investigate how the fraction of sunglasses images impacts the accuracy. First, we downscale the dataset to 500K by randomly selecting the no-sunglasses images from 38,802 identities that are detected having sunglasses images. In this scale, the detected wearing-sunglasses images occupy roughly 23% of the dataset. We then train the models with varying proportions of wearing-sunglasses images by replacing the no-sunglasses images with wearing-sunglasses images. These models are tested on ND-Sunglasses to obtain the FPIR trend w.r.t. different proportions. The left plot in Figure 9 indicates a clear decreasing FPIR when increasing the fraction of sunglasses images. From 0% to 23%, the FPIR has a 42.87% decrease for CF and all CM probe images are correctly identified.
To efficiently add new wearing-sunglasses images to the dataset, we choose to use a face synthesizing model, Vec2Face [34]. It can effectively generate the images of the same identity with sunglasses, as shown in Figure 10. We then conduct the same experiments to evaluate if the synthesized wearing-sunglasses images can achieve the similar effect to the real ones. Note that, we uniformly generate images for all 38,082 identities. The right plot in Figure 9 indicates that using synthesized images can achieve better performance than real ones. The FPIR has a 57.17% decrease for CF and all CM probe images are correctly identified. This showcases the possible accuracy improvement in the full-size WebFace4M by adding synthesized wearing-sunglasses images in the future work. It is worth noting that all the trained models have close average accuracy on five standard test sets [15, 38, 39, 23, 27], as shown in Table VI.
@@ V Conclusions and Future Work
Our research also demonstrates that increasing the proportion of sunglasses-wearing images in the gallery can partially mitigate the accuracy loss caused by sunglasses in probe images. Specifically, adding sunglasses to all gallery images recovered 35.79% of lost accuracy for females and 37.63% for males with ArcFace.
We show that the accuracy degradation caused by wearing sunglasses is due to the under-representation of wearing-sunglasses images. Moreover, we show that using a face synthesizing model to uniformly add wearing-sunglasses images to the dataset can reduce the error rate. Future work is to deploy the same strategy on a full-size training set.

### v2

@@ IV-B Effects of Sunglasses + Blur / Resolution
TABLE II: Wasserstein distance for shift from baseline caused by effects in the probe images. Distributions in Figure 7. Face matcher is AdaFace.
@@
TABLE III: FPIR for one-to-many identification experiments under various effects in the probe images. As an example, with sunglasses + blur in the probe images, 4.522% of Caucasian females probe images had a rank-one false positive identification with AdaFace.
Figure 6 shows example probe images for six conditions: 1) Original high-quality image, 2) Sunglasses added to original, 3) $\sigma=4.6$ Gaussian blur added to original, 4) Original resized to 37x37, 5) Sunglasses and blur, and 6) Sunglasses and reduced resolution. Note that adding sunglasses to an original image may subjectively appear to result in a high-quality image, even though accuracy is degraded similar to noticeable blur or low resolution. Table I shows that when combining blur and low resolution with sunglasses, the d-prime values are almost doubled between mated distributions and almost tripled between non-mated distributions, compared with the single factor, for both matchers and both demographic groups. For example, in the AdaFace results for mated females, the d-prime values are 2.6694 when only adding sunglasses to the probes and 5.783 when combining sunglasses and blur in the probes.
@@ IV-C Analysis of Effect on Model Performance
To quantitatively analyze the FPIs caused by each condition, we report the Wasserstein Distance of the score difference distribution and the FPIR. For a given probe image, a negative score difference indicates an FPI, as the highest similarity value between the probe image and an image from a different person is higher than the top matched image from the same identity. Thus, the distribution of score differences provides a measure of the FPI potential. The distributions of these differences are shown in Figure 7. A general conclusion is that each of the probe quality factors – wearing sunglasses, increased blur, lower resolution – causes the (mated - non-mated) score distribution to shift to lower values, representing increased chances for FPI. The greatest shifts occur when probe images have combined blur and sunglasses, or combined lower-resolution and sunglasses.
@@
We also calculated the FPIR for each probe quality condition, using both ArcFace and AdaFace. These results in Table III show that both ArcFace and AdaFace have zero FPIR for the original and 37x37 resolution probes. AdaFace experiences an FPIR of 0.522% for sunglasses-added probes and an FPIR of 0.174% for blurred probes. But the FPIR increases sharply when sunglasses are combined with either blur or reduced resolution. For example, sunglasses + blur results in FPIR of 3.652% for females for ArcFace, and 4.522% for AdaFace. The absolute level of FPIR varies with the size of a dataset, with a larger number of people enrolled in the gallery causing an increased FPIR [24]. So the overall low FPIR in these results does not project what can be experienced for scenarios with larger galleries. However, the important general point is that multiple degraded quality factors in the probe image causes a dramatic increase in the FPIR. Also, both matchers consistently show higher FPIR for females than for males. This gender-based difference illustrates another general lesson for one-to-many identification accuracy.
@@ IV-D Strategy for Recovering Accuracy Loss
Previous sections show that a probe image with sunglasses matched against gallery images without sunglasses results in a substantial decrease in accuracy. Primarily, the occlusion of facial features by sunglasses in the probe image obscures vital identifying information. But also, comparing images with and without sunglasses presents a mismatch in facial features, which is potentially a different factor affecting accuracy. One could imagine an approach to remove sunglasses from the probe image, or an approach to add sunglasses to the gallery images. Removing sunglasses from the probe could perhaps require less computation, as it involves the one image, but it requires hallucinating face feature content that is unseen. Adding sunglasses to the gallery images, on the other hand, involves adding a plausible instance of a specific style of occlusion to each image.
@@
Table IV shows the d-prime values of rank-one similarity distributions for probe images with sunglasses under different gallery conditions. When sunglasses are randomly added to a single image per identity in the gallery, with an average of 11.15 images per identity, the results show d-prime values of 0.5344 (mated) and 0.7142 (non-mated) for females, and 0.6911 (mated) and 0.891 (non-mated) for males. The d-prime values become more pronounced when sunglasses are added to all gallery images: 1.3203 (mated) and 1.6576 (non-mated) for females, and 1.619 (mated) and 1.8357 (non-mated) for males. Notably, in both scenarios, the improvement in accuracy is more substantial for males than for females. The results demonstrate that under both gallery conditions, accuracy improves. However, adding sunglasses to all gallery images recovers more accuracy. The value of adding sunglasses to the gallery images to ameliorate the accuracy degradation is more obvious in the (mated - non-mated) distributions shown in Figure 8.
@@ IV-E Effect of Sunglasses Training Samples in Accuracy
Accuracy lost by adding sunglasses to face can possibly be explained by the under-representation of wearing-sunglasses images in the training set. To explore this, we start with a widely recognized training set, WebFace4M [41], which has 4M images from 200K identities.
WebFace4M images do not come with meta-data for sunglasses. And we did not find a suitable sunglasses detector already available. Therefore, we developed a sunglasses detector to estimate the number of wearing-sunglasses samples in WebFace4M. Specifically, we first use a facial attribute classifier [37] to get the wearing-eyeglasses images in WebFace4M. Part of these images are then manually separated to three classes: no-eyeglasses, eyeglasses, and sunglasses, for the next training step. This results in 8,500 dark sunglasses images, 8,200 eyeglasses images, and 8,200 no-eyeglasses images. This dataset is subsequently partitioned into training, validation, and test allocations of 70%, 15%, and 15%, respectively. We then fine-tuned a ResNet101 [12], pre-trained on ImageNet, using this dataset. This sunglasses detector achieves 93.3% accuracy on the test set. After running this detector on WebFace4M, 115,818 (3.8%) images are predicted as having dark sunglasses. The strong under-representation of sunglasses in the training set motivates increasing the representation of wearing-sunglasses images in the training set as a possible way to improve accuracy.
@@
However, adding sunglasses images to the dataset is non-trivial, so we investigate how the fraction of sunglasses images impacts the accuracy. First, we downscale the dataset to 500K by randomly selecting the no-sunglasses images from 38,802 identities that are detected as having sunglasses images. In this scale, the detected wearing-sunglasses images occupy roughly 23% of the dataset. We then train the models with varying proportions of wearing-sunglasses images by replacing the no-sunglasses images with wearing-sunglasses images. These models are tested on ND-Sunglasses to obtain the FPIR trend. The left plot in Figure 9 indicates a clear decreasing FPIR when increasing the fraction of sunglasses images. From 0% to 23%, the FPIR has a 42.87% decrease for CF and all CM probe images are correctly identified.
To efficiently add new wearing-sunglasses images to the dataset, we choose to use a face generative model, Vec2Face [35]. It can effectively generate images of the same identity with sunglasses, as shown in Figure 10. We then conduct the same experiments to evaluate if the synthesized wearing-sunglasses images can achieve the similar effect to the real ones. Note that we uniformly generate images for all 38,082 identities. The right plot in Figure 9 indicates that using synthesized images can achieve better performance than real ones. The FPIR has a 57.17% decrease for CF and all CM probe images are correctly identified. This showcases the possible accuracy improvement in the full-size WebFace4M by adding synthesized wearing-sunglasses images in the future work. It is worth noting that all the trained models have close average accuracy on five standard test sets [15, 39, 40, 23, 27], as shown in Table VI.
@@ V Conclusions and Future Work
Our research also demonstrates that generating sunglasses-wearing versions of gallery images can partially mitigate the accuracy loss caused by sunglasses in probe images. Specifically, adding sunglasses to all gallery images recovered 35.79% of lost accuracy for females and 37.63% for males with ArcFace.
We show that the accuracy degradation for probe images with sunglasses can also be addressed by fixing the under-representation of wearing-sunglasses images in the training data. We show that using a face generative model to uniformly add wearing-sunglasses images to the training set can reduce the FPIR. Future work is to explore versions of this strategy on a full-size training set.
@@ VI Ethical Impact Statement
Supplementary Material
The experiments in the paper primarily utilize the AdaFace matcher. This supplementary material presents the same experimental results conducted with the ArcFace recognition matcher, thereby validating the conclusions of the paper. Specifically, this supplementary material provides:
  • Figure 11 shows some examples of different types of synthetic sunglasses added by the FaceLab app. This app will add different types of realistic sunglasses based on the images.
  • Figure 12 demonstrates the high identity-preservation of sunglasses augmentation produced by the FaceLab app. The minimal differences and high similarity between the original and augmented images validate the use of FaceLab app for sunglasses addition in our research.
  • Figure 13 shows the FaceLab app’s modifications are localized exclusively to the sunglasses’ region.
  • Figure 17 shows shows the probe with sunglasses + blur, the gallery image that is the highest similarity of the right identity, and the gallery image that matched to it as the rank-one similarity but the wrong person.
  • Figure 18 shows a comparison of mated and non-mated distributions for original probe images and various conditions of degraded probes using the AdaFace matcher.
  • Figure 19 shows a comparison of mated and non-mated distributions for original probe images and various conditions of degraded probes using the ArcFace matcher.
  • Figure 20 shows score difference (mated - non-mated) distributions for original probe images and different conditions of degraded probes using the ArcFace matcher.
  • Figure 21 compares mated and non-mated distributions for sunglasses-added probe images under different gallery degradation conditions using the ArcFace matcher.
  • Figure 22 compares mated and non-mated distributions for sunglasses-added probe images under different gallery degradation conditions using the AdaFace matcher.
  • Figure 23 compares difference (mated - non-mated) distributions for sunglasses-added probe images under different gallery degradation conditions using the ArcFace matcher.
  • Table VII shows the Wasserstein distances of the differences in (mated - non-mated) distributions in Figure 19, under various effects to the probe images.
Figure 11: Examples of different types of synthetic sunglasses added by FaceLab app.
Figure 12: This study illustrates a small experiment comprising three conditions: a) the original image, b) the image with physical sunglasses applied, and c) the image with sunglasses using the FaceLab app. Absolute differences between these images were calculated, along with cosine similarities between the original and each modified version. Results revealed cosine similarities of 0.773 for physical sunglasses and 0.819 for FaceLab-applied sunglasses. The higher similarity of FaceLab-augmented images to the original supports the validity of using the FaceLab application for sunglasses augmentation in our research methodology.
Figure 13: The figure comprises three columns: the original image, the FaceLab app-modified image with added sunglasses, and the absolute difference between the two. This visualization demonstrates that the FaceLab app’s modifications are localized exclusively to the sunglasses’ region.
(a) Sunglasses + blur Probe
(b) Mated Match Gallery
(c) Non-Mated Match Gallery
(d) (a) Caucasian Females
(e) (b) Caucasian Males
Figure 17: This figure shows the probe with sunglasses + blur, the gallery image that is the highest similarity of the right / same identity, and the gallery image that matched to it as the rank-one similarity but the wrong / different person.
Figure 18: Comparison mated and non-mated distributions for original probe images and different conditions of degraded probes (top to bottom): 1) sunglasses, 2) blur of $\sigma=4.6$, 3) 37x37 resolution, 4) sunglasses + blur, 5) sunglasses + 37x37 resolution. The face matcher is AdaFace.
Figure 19: Comparison mated and non-mated distributions for original probe images and different conditions of degraded probes (top to bottom): 1) sunglasses, 2) blur of $\sigma=4.6$, 3) 37x37 resolution, 4) sunglasses + blur, 5) sunglasses + 37x37 resolution. The face matcher is ArcFace.
Figure 20: Score difference (mated - non-mated) distributions for original probe images and different conditions of degraded probes (top to bottom): 1) sunglasses, 2) blur of $\sigma=4.6$, 3) 37x37 resolution, 4) sunglasses + blur, 5) sunglasses + 37x37 resolution. The face matcher is ArcFace.
Figure 21: Comparing mated and non-mated distributions for sunglasses-added probe images under different gallery degradation conditions: (a) sunglasses randomly added to one image per identity in the gallery, (b) sunglasses added to all gallery images. Distributions are shown separately for male and female subjects. It used the ArcFace recognition matcher.
Figure 22: Comparing mated and non-mated distributions for sunglasses-added probe images under different gallery degradation conditions: (a) sunglasses randomly added to one image per identity in the gallery, (b) sunglasses added to all gallery images. Distributions are shown separately for male and female subjects. It used the AdaFace recognition matcher.
Figure 23: Similarity difference (mated - non-mated) distributions. Comparing with the baseline, adding sunglasses to probe images causes a large accuracy loss. To counter this effect, we add sunglasses to one (all) image(s) per identity in the gallery to achieve good accuracy recovery, as shown in Top (Bottom) row. It used the ArcFace recognition matcher.
Effects added to probe images Demographic Probe vs. original gallery
Sunglasses CF 0.197
CM 0.194
Gaussian blur ($\sigma=4.6$) CF 0.190
CM 0.187
Low-resolution (37x37) CF 0.199
CM 0.191
Sunglasses and Gaussian blur CF 0.366
CM 0.354
Sunglasses and low-resolution CF 0.366
CM 0.334
TABLE VII: Wasserstein distances of the differences in (mated - non-mated) distributions in Figure 19, under various effects to the probe images.